# Rabble-Rousers in the New King’s Court: Algorithmic Effects on Account Visibility in Pre-X Twitter

### v1

@@ Abstract
Algorithmic effects on social media platforms have come under recent scrutiny, with several works reporting that right-leaning accounts tend to receive more exposure. In this paper, we expand upon this body of work using data collected from user feeds after Twitter’s change of ownership but before its re-branding to X. We replicate findings from prior work regarding the increased exposure of right-leaning accounts to wider audiences in algorithmically curated compared to reverse-chronological feeds, and, crucially, we further unpack this effect to understand what correlated (and did not correlate) with these differences. Our results reveal that right-leaning accounts benefited not necessarily due to their political affiliation, but possibly because they behaved in ways associated with algorithmic rewards; namely, posting more agitating content and receiving attention from the platform’s owner, Elon Musk, who was the most central network account. We also demonstrate that legacy-verified accounts, like businesses and government officials, received less exposure in the algorithmic feed compared to non-verified or Twitter Blue-verified accounts. We discuss implications of these findings for the intersection between behavioral incentives for algorithmic reach and online trust and safety.
@@ Introduction
The role of algorithms in amplifying divisive and problematic content has been a question of both societal (pariser_filter_2011) and academic concern (ribeiro_auditing_2020). Recently, debates have sparked around whether recommendation algorithms disproportionately amplify or suppress content from specific political camps. For example, the Federal Trade Commission under the second Trump administration launched an investigation into alleged “censorship” of conservative voices, [^2] https://archive.ph/xcShX while other reports actually suggest more amplification of right-leaning accounts (graham_computational_2024). Yet, others, including former Twitter scientists (messing_what_2023), have argued that algorithmic effects are deceptively difficult to measure because of algorithms’ inherent reliance on user preferences that shape what these algorithms learn (ribeiro_amplification_2023).
The challenge and urgency of this debate has prompted research aimed at isolating and characterizing algorithmic effects in political contexts. Much of this work has relied on the use of automated accounts, or “bots” to capture algorithmically recommended content (ye_auditing_2025; Duskin2026-new; Bandy2021-lg). Recent work has also incorporated “counterfactual bots” to control for baseline user behavior (hosseinmardi_causally_2024). That is, a pair of bots, one of which is instructed to behave like a real user and the other instructed to randomly follow algorithmic recommendations, can be compared to estimate content amplified beyond baseline user preferences.
@@
We use data from milli_engagement_2025’s study on algorithmic elevation of emotional and outgroup-derogating content, who collected simultaneous algorithmic and chronological Twitter feeds from participants recruited from an online panel. Building upon their work, we focus our analyses on account level amplification, rather than the tweet level, so that we can also capture account characteristics like social proximity to the network center.
@@
Among the influential recent work sparking interest in algorithmic audits is ribeiro_auditing_2020, which traced video recommendations on YouTube to chart radicalization pathways from pseudo-intellectualism to the alt-right. These types of studies proliferated on YouTube (haroon_auditing_2023; ibrahim_youtube_2023; hosseinmardi_causally_2024) primarily because its API used to offer endpoints that enabled this kind of study, something which is no longer the case. Since then, multiple other studies have emulated similar methods by deploying automated accounts that simulate users (aka, sockpuppets) and observing the content recommended to them. This includes studies on Twitter’s algorithmic timeline, and geolocation-based SERP audits for COVID-19 misinformation on YouTube (jung_algorithmic_2025), among others. The benefit of these automated audits is that they are not obfuscated by user activity, allowing for the study of algorithmic baselines or priors.
@@
One of the largest such studies was conducted by the Twitter team in collaboration with academics (huszar_algorithmic_2022). This work, which randomized $\sim$2M Twitter users into a reverse-chronological feed, found that algorithmic amplification favored right-leaning politician accounts and news sources compared to left-leaning ones, although it did not find evidence of amplification for users belonging to extreme groups. Most similar to our work, milli_engagement_2025, whose data we use in this study, collected both the engagement and reverse-chronological Twitter feeds of the same users at the same time. They found that the algorithmic feed featured more emotional and outgroup-derogating content, although this was not necessarily content that users reported preferring. We expand beyond this content-centered analysis to consider how user-level characteristics and interaction patterns are associated with visibility within the algorithmic feed.
@@
Despite notable contributions and a growing body of prior work, we still lack an adequate understanding of the kinds of accounts that algorithmic manifestations benefit. Indeed, few works focus at the account level (ye_auditing_2025). We argue for addressing this gap because 1) it allows us to capture network effects, specifically, the proximity of accounts to the most central network node, that are inherently built into algorithmic recommendations and 2) it allows us to better understand social media incentives in gaining influence, which are paramount for regulators and legislators. Moreover, although some works offer rich descriptions of what kind of content may benefit algorithmically (milli_engagement_2025), we can paint a fuller picture with more research that considers previously unexamined dimensions.
@@ Methods
We use a dataset made available by milli_engagement_2025, who collected both the reverse-chronological and “For You” (henceforth, “engagement” as per the original paper) Twitter feeds of 806 US residents recruited on the CloudResearch Connect panel platform between February 11th-27th, 2023. Of note is that, despite the somewhat smaller sample and limited time span, these data were collected one month before Twitter released code for its recommendation algorithm on March 31st 2023, [^4] https://github.com/twitter/the-algorithm making it highly likely that our observations are driven by the same or very similar version of that algorithm. This allows for a unique juxtaposition between what the algorithm was built to do versus what it did do, namely, the stated purpose of recommending more interesting content versus the potential for amplifying more problematic content or over-centralizing recommendations among a few important accounts.
Across both feeds and all participants, this dataset captures 205k potential exposures to 171k unique tweets authored or retweeted by 63.4k unique accounts. We direct the reader to the original paper for detailed participant demographics. However, given the politically-adjacent focus of this paper, we clarify that the dataset is heavily skewed with 76.7% of participants identifying as left-leaning and 23.3% as right-leaning, suggesting that the sample may not be nationally representative.555milli_engagement_2025 also report that the Twitter user population skewed Democrat in a 2020 ANES study of a nationally representative sample, but not to the degree of this dataset. Therefore, we make balancing adjustments where necessary. The dataset does not contain any promotional or ad tweets.
@@
To address the political skew of the sample, we match the minority right-leaning participants to a subset of left-leaning ones, using the demographic variables that milli_engagement_2025 obtained. [^6] See https://github.com/smilli/twitter/blob/main/DATA.md for potential responses. We form participant vectors consisting of the categorical variables race, gender, and reason for using Twitter (e.g., entertainment, to stay informed, etc.), and the ordinal variables education level, age group, and annual household income. Categorical variables are one-hot encoded, and all (resulting) variables are equally weighted. We compute pairwise cosine distances between all right-leaning to left-leaning participants, and perform nearest-neighbor matching without replacement (such that each right-leaning user is matched to exactly one unique left-leaning user). Whenever we refer to a “balanced network” henceforth, we mean a network based on this matched set of participants. We show the (unweighted) balanced networks based on the reverse-chronological and engagement feeds in Figure 1. Results presented in the next section are largely consistent with analyses utilizing the full sample. We demonstrate the success of our matching procedure in the Appendix.
@@
We begin with a description of the two unweighted feed configurations in Table 1. To compute partisan assortativity, we project the networks such that an edge forms between two participants if they are exposed to the same account. To compute (in-)degree centralization, we implement borgatti_network_1997’s method for bipartite graphs that computes the theoretical maximum degree centralization by accounting for the cardinality of different vertex sets instead of relying on a unipartite star graph as follows:
@@
Although there was a larger set of exposed accounts (and, by extension, $P\rightarrow A$ edges) in the reverse-chronological feed, this feed also showed higher partisan assortativity, meaning that there was more partisan homogeneity in the accounts that participants were exposed to. This is in line with several recent works that suggest partisan sorting may mostly arise due to user preference rather than algorithmic recommendations (chouaki_what_2024; duskin_echo_2024_new; robertson_users_2023). However, we also notice much higher centralization in the engagement feed, suggesting that exposure was more concentrated among a few important accounts. We explore this finding next. The patterns we observe are identical when preserving the entire participant sample (N = 806).
@@
To better illustrate the kinds of accounts that featured most prominently in the chronological and engagement feed, we show the top-10 for each feed in Table 2. This also demonstrates the ability of our method for determining political leaning to distinguish between well-known left- and right-leaning accounts on Twitter. The table largely reflect the patterns in Figure 2; left-leaning accounts reached more users and were more important in the chronological feed, whereas the engagement feed featured right-leaning accounts more heavily. We also notice a substantial difference in centrality distributions, with Elon Musk’s account (the most influential in both feeds) becoming much more centralized in the engagement feed relative to the second most-central account (0.52 and 0.20, respectively) when compared to the chronological feed (0.27 and 0.21, respectively). Thus, Elon Musk was likely the main driver of the substantially higher engagement feed centralization we report in Table 1.
@@
For each tweet, we use an LLM (Gemini 2.0 Flash) to annotate whether it was political or not. We test the model’s performance against a subset of 30.6k tweets annotated by the participants in milli_engagement_2025 (after removing URL-only tweets); the prompt we use for the LLM is adapted from the question shown to participants, which also captures social issues. As we show in Figure 5, the model performs well and trends towards consensus; that is, performance is increased when considering tweets annotated by more participants that enable us to take the majority label. For each account, we then compute the percentage of their in-sample tweets that were political.
@@
We also use the same model to annotate whether a tweet was agitating/conflict-inducing (see Appendix for exact prompt). Although milli_engagement_2025’s dataset contains ratings for sentiment (e.g., anxiety, happiness) and anger expressed towards the left or right in a subset of tweets, analyses on GPT-4 labels provided in the original dataset reveal that LLM agreement with human annotators on these categories tends to be lower. Thus, we opt for this custom agitating annotation, which we define as a tweet that stirs controversy or conflict. We argue that this can capture more subtle ways of inducing negative responses and is more closely tied to the contents of the tweet itself, compared to commonly-used metrics like emotionality, toxicity, or identity attack which may also be largely dependent on the perceiver’s identity (aroyo_crowdsourcing_2019; goyal_is_2022). For example, a tweet like “Politician X is trying to destroy America” could agitate both supporters and opponents of that politician alike; however, this tweet is neither toxic, nor does it directly attack any kind of identity.
@@ Discussion
Our findings shed light on how Twitter’s recommendation algorithm prior to its re-branding to X may have affected the visibility of different accounts. While we replicate the finding reported by several works (graham_computational_2024; huszar_algorithmic_2022; Duskin2026-new; ye_auditing_2025) that right-leaning accounts benefited from algorithmic curation, upon closer inspection, we find that this was likely due to them acting in ways that correlated with algorithmic rewards. Namely, right-leaning accounts posted more agitating content and were closer to Elon Musk, both of which were associated with more algorithmic exposure.
At the same time, contrary to previous work (huszar_algorithmic_2022), we find that accounts of government officials (and businesses, which are also legacy-verified) lost visibility in the algorithmic configuration. There could be several reasons behind this, including different timelines, operationalization of government accounts, and the fact that huszar_algorithmic_2022 did not only focus on the US context.
@@
Additionally, our findings highlight a troublesome implication. In the pursuit of reaching wider audiences, users may be incentivized to stir controversy or vie for engagement with the platform’s owner, creating an environment of elevated agitation and inadvertent “permissible reach”. Indeed, some work has demonstrated an uptick in problematic behavior such as hate speech and automated activity following Musk’s acquisition of the platform (hickey_auditing_2023; hickey_x_2025). This is antithetical to the platform’s ostensible role as a “digital town square”, especially in the wake of cuts to the Trust and Safety team, which would perhaps be best placed to monitor and address such issues (moran_end_2025_new).
Moreover, we add to literature demonstrating political differences in the adoption of problematic behaviors (mosleh_differences_2024), which possibly explains the different rates at which different political groups are subjected to content moderation (haimson_disproportionate_2021; renault_republicans_2025). Thus, our work helps to further contextualize research on the types of content that achieve more visibility (galeazzi_revealing_2026), as well as discussions around why accusations of bias may ignore crucial context behind these moderation decisions. [^10] https://www.techpolicy.press/scientists-respond-to-ftc-inquiry-into-tech-censorship/
@@
Although we perform comparisons between feeds for the same users, our findings are not necessarily causal. Without systematic randomization of users into different configurations (guess_how_2023; huszar_algorithmic_2022) or counterfactual behaviors (hosseinmardi_causally_2024) it is difficult to disentangle algorithmic amplification from overall user preferences. However, given that we are essentially comparing “counterfactual feeds”, these findings are a good description of how user experience differed in algorithmic versus reverse-chronological configurations.
@@ Acknowledgments
This work has been supported by the University of Washington’s Center for an Informed Public, the John S. and James L. Knight Foundation (G-2019-58788), and the William and Flora Hewlett Foundation (2023-02789). The authors thank Akhil Chennamsetty and Darren Linvill for providing the Musk interaction data, and Martin Saveski for his helpful feedback on earlier drafts of this paper.

### v2

@@ Abstract
Algorithmic effects on social media platforms have come under recent scrutiny, with several studies reporting that right-leaning accounts tend to receive more exposure. In this paper, we expand upon this body of work using data collected from user feeds after Twitter’s change of ownership but before its re-branding to X. We replicate findings from prior work regarding the increased exposure of right-leaning accounts to wider audiences in algorithmically curated compared to reverse-chronological feeds, and, crucially, we further unpack this effect to illuminate what correlated (and did not correlate) with these differences. Our results reveal that right-leaning accounts benefited not necessarily due to their political affiliation, but likely because they behaved in ways associated with algorithmic rewards; namely, posting more agitating content and receiving attention from the platform’s owner, Elon Musk, who was the most central network account. We also demonstrate that legacy-verified accounts, like businesses and government officials, received less exposure in the algorithmic feed compared to non-verified or Twitter Blue-verified accounts. We discuss implications of these findings for the intersection between behavioral incentives for algorithmic reach and the health of online discourse.
@@ Introduction
The role of algorithms in amplifying divisive and problematic content is a question of both societal (Pariser 2011) and academic concern (Ribeiro et al. 2020). Recently, debates have sparked around whether recommendation algorithms disproportionately amplify or suppress content from specific political camps. For example, the Federal Trade Commission under the second Trump administration launched an investigation into alleged “censorship” of conservative voices, [^2] https://archive.ph/xcShX while other reports suggest right-leaning accounts actually receive outsized amplification (Graham and Andrejevic 2024). Yet, others, including former Twitter employees (Messing 2023), have argued that algorithmic effects are deceptively difficult to measure because of algorithms’ inherent reliance on user preferences that shape what these algorithms learn (Ribeiro et al. 2023).
The challenge and urgency of this debate has prompted research aimed at isolating and characterizing algorithmic effects in political contexts. Much of this work has relied on the use of automated accounts, or “bots” to capture algorithmically recommended content (Ye et al. 2025; Duskin et al. 2025; Bandy and Diakopoulos 2021). Recent work has also incorporated “counterfactual bots” to control for baseline user behavior (Hosseinmardi et al. 2024). That is, a pair of bots, one of which is instructed to behave like a real user and the other instructed to randomly follow algorithmic recommendations, can be compared to estimate content amplified beyond baseline user preferences.
@@
We use data from Milli et al. (2025)’s study on algorithmic elevation of emotional and outgroup-derogating content, who collected simultaneous algorithmic and chronological Twitter feeds from participants recruited from an online panel. Building upon their work, we focus our analyses on account level amplification, rather than the tweet level, so that we can also capture account characteristics like social proximity to the network center.
@@
Among the influential recent work sparking interest in algorithmic audits is Ribeiro et al. (2020), which traced video recommendations on YouTube to chart radicalization pathways from pseudo-intellectualism to the alt-right. These types of studies proliferated on YouTube (Haroon et al. 2023; Ibrahim et al. 2023; Hosseinmardi et al. 2024) primarily because its API used to offer endpoints that enabled this kind of study, something which is no longer the case. Since then, multiple other studies have emulated similar methods by deploying automated accounts that simulate users (aka, sockpuppets) and observing the content recommended to them. This includes studies on Twitter’s algorithmic timeline, and geolocation-based SERP audits for COVID-19 misinformation on YouTube (Jung et al. 2025), among others. The benefit of these automated audits is that they are not obfuscated by user activity, allowing for the study of algorithmic baselines or priors.
@@
One of the largest such studies was conducted by the Twitter team in collaboration with academics (Huszár et al. 2022). This work, which randomized $\sim$2M Twitter users into a reverse-chronological feed, found that algorithmic amplification favored right-leaning politician accounts and news sources compared to left-leaning ones, although it did not find evidence of amplification for users belonging to extreme groups. Most similar to our work, Milli et al. (2025), whose data we use in this study, collected both the engagement and reverse-chronological Twitter feeds of the same users at the same time. They found that the algorithmic feed featured more emotional and outgroup-derogating content, although this was not necessarily content that users reported preferring. Here, we expand beyond this content-centered analysis to consider how user-level characteristics and interaction patterns are associated with visibility within the algorithmic feed.
@@
Despite notable contributions and a growing body of prior work, we still lack an adequate understanding of the kinds of accounts that algorithmic manifestations benefit. Indeed, few works focus at the account level (Ye et al. 2025). We argue for addressing this gap because 1) it allows us to capture network effects, specifically, the proximity of accounts to the most central network node, that are inherently built into algorithmic recommendations and 2) it allows us to better understand social media incentives in gaining influence, which are paramount for regulators and legislators. Moreover, although some works offer rich descriptions of what kind of content may benefit algorithmically (Milli et al. 2025), we can paint a fuller picture with more research that considers previously unexamined dimensions.
@@ Methods
We use a dataset made available by Milli et al. (2025), who collected both the reverse-chronological and “For You” (henceforth, “engagement” as per the original paper) Twitter feeds of 806 US residents recruited on the CloudResearch Connect panel platform between February 11th-27th, 2023. Though the collection window was short and the resulting dataset is relatively small, these data offer insight into a particular salient moment in the platform’s history: they were collected (1) as the platform was changing identity in response to new ownership and (2) one month before Twitter released code for its recommendation algorithm on March 31st 2023, [^4] https://github.com/twitter/the-algorithm making it highly likely that our observations are driven by the same or very similar version of that algorithm. This allows for a unique juxtaposition between what the algorithm was built to do versus what it did do, namely, the stated purpose of recommending more interesting content versus the potential for amplifying more problematic content or over-centralizing recommendations among a few important accounts.
Across both feeds and all participants, this dataset captures 205k potential exposures to 171k unique tweets authored or retweeted by 63.4k unique accounts. We direct the reader to the original paper for detailed participant demographics. However, given the politically-adjacent focus of this paper, we clarify that the dataset is heavily skewed with 76.7% of participants identifying as left-leaning and 23.3% as right-leaning, suggesting that the sample may not be nationally representative. [^5] Milli et al. (2025) also report that the Twitter user population skewed Democrat in a 2020 ANES study of a nationally representative sample, but not to the degree of this dataset. Therefore, we make balancing adjustments where necessary. The dataset does not contain any promotional or ad tweets.
@@
To address the political skew of the sample, we match the minority right-leaning participants to a subset of left-leaning ones, using the demographic variables that Milli et al. (2025) obtained. [^6] See https://github.com/smilli/twitter/blob/main/DATA.md for potential responses. We form participant vectors consisting of the categorical variables race, gender, and reason for using Twitter (e.g., entertainment, to stay informed, etc.), and the ordinal variables education level, age group, and annual household income. Categorical variables are one-hot encoded, and all (resulting) variables are equally weighted. We compute pairwise cosine distances between all right-leaning to left-leaning participants, and perform nearest-neighbor matching without replacement (such that each right-leaning user is matched to exactly one unique left-leaning user). Whenever we refer to a “balanced network” henceforth, we mean a network based on this matched set of participants. We show the (unweighted) balanced networks based on the reverse-chronological and engagement feeds in Figure 1. Results presented in the next section are largely consistent with analyses utilizing the full sample. We demonstrate the success of our matching procedure in the Appendix.
@@
We begin with a description of the two unweighted feed configurations in Table 1. To compute partisan assortativity, we project the networks such that an edge forms between two participants if they are exposed to the same account. To compute (in-)degree centralization, we implement Borgatti and Everett (1997)’s method for bipartite graphs that computes the theoretical maximum degree centralization by accounting for the cardinality of different vertex sets instead of relying on a unipartite star graph as follows:
@@
Although there was a larger set of exposed accounts (and, by extension, $P\rightarrow A$ edges) in the reverse-chronological feed, this feed also showed higher partisan assortativity, meaning that there was more partisan homogeneity in the accounts that participants were exposed to. This is in line with several recent works that suggest partisan sorting may mostly arise due to user preference rather than algorithmic recommendations (Chouaki et al. 2024; Duskin et al. 2024; Robertson et al. 2023). However, we also notice much higher centralization in the engagement feed, suggesting that exposure was more concentrated among a few important accounts. We explore this finding next. The patterns we observe are identical when preserving the entire participant sample (N = 806).
@@
To better illustrate the kinds of accounts that featured most prominently in the chronological and engagement feed, we show the top-10 for each feed in Table 2. This also demonstrates the ability of our method for determining political leaning to distinguish between well-known left- and right-leaning accounts on Twitter. The table largely reflects the patterns in Figure 2; left-leaning accounts reached more users and were more important in the chronological feed, whereas the engagement feed featured right-leaning accounts more heavily. We also notice a substantial difference in centrality distributions, with Elon Musk’s account (the most influential in both feeds) becoming much more centralized in the engagement feed relative to the second most-central account (0.52 and 0.20, respectively) when compared to the chronological feed (0.27 and 0.21, respectively). Thus, Elon Musk was likely the main driver of the substantially higher engagement feed centralization we report in Table 1.
@@
For each tweet, we use an LLM (Gemini 2.0 Flash) to annotate whether it was political or not. We test the model’s performance against a subset of 30.6k tweets annotated by the participants in Milli et al. (2025) (after removing URL-only tweets); the prompt we use for the LLM is adapted from the question shown to participants, which also captures social issues. As we show in Figure 5, the model performs well and trends towards consensus; that is, performance is increased when considering tweets annotated by more participants that enable us to take the majority label. For each account, we then compute the percentage of their in-sample tweets that were political.
@@
We also use the same model to annotate whether a tweet was agitating/conflict-inducing (see Appendix for exact prompt). Although Milli et al. (2025)’s dataset contains ratings for sentiment (e.g., anxiety, happiness) and anger expressed towards the left or right in a subset of tweets, analyses on GPT-4 labels provided in the original dataset reveal that LLM agreement with human annotators on these categories tends to be lower. Thus, we opt for this custom agitating annotation, which we define as a tweet that stirs controversy or conflict. We argue that this can capture more subtle ways of inducing negative responses and is more closely tied to the contents of the tweet itself, compared to commonly-used metrics like emotionality, toxicity, or identity attack which may also be largely dependent on the perceiver’s identity (Aroyo et al. 2019; Goyal et al. 2022). For example, a tweet like “Politician X is trying to destroy America” could agitate both supporters and opponents of that politician alike; however, this tweet is neither toxic, nor does it directly attack any kind of identity.
@@ Discussion
Our findings shed light on how Twitter’s recommendation algorithm prior to its re-branding to X may have affected the visibility of different accounts. While we replicate the finding reported by several works (Graham and Andrejevic 2024; Huszár et al. 2022; Duskin et al. 2025; Ye et al. 2025) that right-leaning accounts benefited from algorithmic curation, upon closer inspection, we find that this was likely due to them acting in ways that correlated with algorithmic rewards. Namely, right-leaning accounts posted more agitating content and were closer to Elon Musk, both of which were associated with more algorithmic exposure.
At the same time, contrary to previous work (Huszár et al. 2022), we find that accounts of government officials (and businesses, which are also legacy-verified) lost visibility in the algorithmic configuration. There could be several reasons behind this, including different timelines, operationalization of government accounts, and the fact that Huszár et al. (2022) did not only focus on the US context.
@@
Additionally, our findings highlight a troublesome implication. In the pursuit of reaching wider audiences, users may be incentivized to stir controversy or vie for engagement with the platform’s owner, creating an environment of elevated agitation and inadvertent “permissible reach”. Indeed, some work has demonstrated an uptick in problematic behavior such as hate speech and automated activity following Musk’s acquisition of the platform (Hickey et al. 2023, 2025). This is antithetical to the platform’s ostensible role as a “digital town square”, especially in the wake of cuts to the Trust and Safety team, which would perhaps be best placed to monitor and address such issues (Moran et al. 2025).
Moreover, we add to literature demonstrating political differences in the adoption of problematic behaviors (Mosleh et al. 2024), which possibly explains the different rates at which different political groups are subjected to content moderation (Haimson et al. 2021; Renault et al. 2025). Thus, our work helps to further contextualize research on the types of content that achieve more visibility (Galeazzi et al. 2026), as well as discussions around why accusations of bias may ignore crucial context behind these moderation decisions. [^10] https://www.techpolicy.press/scientists-respond-to-ftc-inquiry-into-tech-censorship/
@@
Although we perform comparisons between feeds for the same users, our findings are not necessarily causal. Without systematic randomization of users into different configurations (Guess et al. 2023; Huszár et al. 2022) or counterfactual behaviors (Hosseinmardi et al. 2024) it is difficult to disentangle algorithmic amplification from overall user preferences. However, given that we are essentially comparing “counterfactual feeds”, these findings are a good description of how user experience differed in algorithmic versus reverse-chronological configurations.
@@ Acknowledgments
This work has been supported by the University of Washington’s Center for an Informed Public, the John S. and James L. Knight Foundation (G-2019-58788), and the William and Flora Hewlett Foundation (2023-02789). The authors thank Akhil Chennamsetty and Darren Linvill for providing the Musk interaction data.